OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, 4>, v = <7, 5> MEDAL!!!!!

OpenStudy (anonymous):

@ganeshie8

OpenStudy (ipwnbunnies):

You can use the dot product derivation. \[u . v = |u||v|\cos \theta\]

OpenStudy (ipwnbunnies):

\[\cos \theta = \frac{u . v}{|u||v|}\]

OpenStudy (anonymous):

i got 1.8 degrees

OpenStudy (anonymous):

did i do it right?

OpenStudy (ipwnbunnies):

Yasss. I also got 1.8 degrees.

OpenStudy (anonymous):

thanks!